Kevin is 16 years old and Vanessa is 2 years old.
How many years will it take until Kevin is only 3 times as old as Vanessa?

Answer by ewatrrr(24785)

x the time it will take until Kevin is 3 times as old as Vanessa
16 + x = 3(2 + x)
16 + x = 6 + 3x
10 = 2x
5 = x
5 years it will take until Kevin is 3 times as old as Vanessa